# The iterated Dirichlet process and applications to Bayesian inference

### v1

@@ 2.1 Random measures and kernels
    $\mu$ is $(\mathcal{T},\mathcal{M})$-measurable.
A probability kernel is a kernel taking values in $M_{1}$. Note that a random measure on $S$ is just a kernel from $\Omega$ to $S$, and a random probability measure on $S$ is just a probability kernel from $\Omega$ to $S$.
It can be shown that the function $\mu\mapsto\mu^{\infty}$ mapping $M_{1}$ to $M_{1}(S^{\infty})$ is measurable. Therefore, if $\mu$ is a random probability measure on $S$, then $\mu^{\infty}$ is a random probability measure on $S^{\infty}$. In fact, the random variable $\mu^{\infty}$ is $\sigma(\mu)$-measurable.
@@
Let $S^{\prime}$ be another complete and separable metric space. Let $\gamma$ be a probability kernel from $T$ to $S$ and $\gamma^{\prime}$ a probability kernel from $T\times S$ to $S^{\prime}$. (We allow the possibility that $T$ is a singleton, in which case $\gamma$ is a probability measure on $S$ and $\gamma^{\prime}$ is a probability kernel from $S$ to $S^{\prime}$.) We write $\gamma\gamma^{\prime}$ to denote the probability kernel from $T$ to $S\times S^{\prime}$ characterized by
$(\gamma\gamma^{\prime})(y,A\times A^{\prime})=\int_{A}\gamma^{\prime}(y,z,A^{\prime})\,\gamma(y,dz),$
where $y\in T$, $z\in S$, $A\in\mathcal{S}$, and $A^{\prime}\in\mathcal{S}^{\prime}$. In particular, this means
@@
2.2 Regular conditional distributions
If $X$ is an $S$-valued random variable and $\mu$ is a probability measure on $S$, then we write $X\sim\mu$, as usual, to mean that $X$ has distribution $\mu$. We also use $\mathcal{L}(X)$ to denote the distribution of $X$, so that $X\sim\mu$ and $\mathcal{L}(X)=\mu$ are synonymous. If $\nu$ is a finite, nonzero measure on $S$, then we write $\mathcal{L}(X)\propto\nu$ to mean that $\mathcal{L}(X)=\nu/\nu(S)$. We will sometimes combine this with differential notation. For example, $d\mathcal{L}(X)\propto f(x)\,\nu(dx)$ means that for some constant $c\in(0,\infty)$, the function $cf$ is the Radon-Nikodym derivative of $\mathcal{L}(X)$ with respect to $\nu$. A naked differential indicates Lebesgue measure, so that in the case $S=\mathbb{R}$, the notation $d\mathcal{L}(X)=f(x)\,dx$ means that $f$ is the density of $X$ with respect to Lebesgue measure. If $\mathcal{L}(X)=\mu$ and $B\in\mathcal{S}$, then we use $\mathcal{L}(X;B)$ to denote $\mu(B)$. In this way, we may sometimes write $\mathcal{L}(X;dx)$ instead of $d\mathcal{L}(X)$.
Let $\mathcal{G}\subseteq\mathcal{F}$ be a $\sigma$-algebra. A regular conditional distribution for $X$ given $\mathcal{G}$ is a random probability measure $\mu$ on $S$ such that $P(X\in A\mid\mathcal{G})=\mu(A)$ a.s. for all $A\in\mathcal{S}$. Since $S$ is a complete and separable metric space, regular conditional distributions exist. In fact, such a $\mu$ exists whenever $X$ takes values in a standard Borel space. Moreover, $\mu$ is unique in the sense that if $\mu$ and $\widetilde{\mu}$ are two such random measures, then $\mu=\widetilde{\mu}$ a.s. That is, with probability one, $\mu(B)=\widetilde{\mu}(B)$ for all $B\in\mathcal{S}$. The random measure $\mu$ may be chosen so that it is $\mathcal{G}$-measurable.
We use the notation $\mathcal{L}(X\mid\mathcal{G})$ to refer to the regular conditional distribution of $X$ given $\mathcal{G}$. If $\mu$ is a random probability measure, the notation ${X\mid\mathcal{G}}\sim\mu$ means $\mathcal{L}(X\mid\mathcal{G})=\mu$ a.s. If $\nu$ is a random measure that is a.s. finite and nonzero, then $\mathcal{L}(X\mid\mathcal{G})\propto\nu$ means ${X\mid\mathcal{G}}\sim\nu/\nu(S)$. As with unconditional distributions, we will combine this with differential and semicolon notation, so that, for instance, $\mathcal{L}({X\mid\mathcal{G}};A)=P(X\in A\mid\mathcal{G})$ a.s.
Let $(T,\mathcal{T})$ be a measurable space and $Y$ a $T$-valued random variable. Then there exists a probability kernel $\mu$ from $T$ to $S$ such that $X\mid\sigma(Y)\sim\mu(Y)$. Moreover, $\mu$ is unique in the sense that if $\mu$ and $\widetilde{\mu}$ are two such probability kernels, then $\mu=\widetilde{\mu}$, $\mathcal{L}(Y)$-a.e. In this case, we will typically omit the $\sigma$, and write, for instance, that $\mathcal{L}(X\mid Y)=\mu(Y)$. Note that for fixed $y\in T$, the probability measure $\mu(y,\cdot)$ is not uniquely determined, since the kernel $\mu$ is only unique $\mathcal{L}(Y)$-a.e. Nonetheless, if a particular $\mu$ has been fixed, we will use the notation $\mathcal{L}(X\mid Y=y)$ to denote the probability measure $\mu(y,\cdot)$.
2.3 Exchangeability
A sequence $\xi=\{\xi_{i}:i\in\mathbb{N}\}$ of $S$-valued random variables is conditionally i.i.d. if there exists a $\sigma$-algebra $\mathcal{G}\subseteq\mathcal{F}$ and a random probability measure $\nu$ on $S$ such that $P(\xi\in A\mid\mathcal{G})=\nu^{\infty}(A)$ a.s., for all $A\in\mathcal{S}^{\infty}$. It can be shown that a sequence $\xi$ is conditionally i.i.d. if and only if there is a random probability measure $\mu$ such that
[…]
and in this case, $\mu$ can be chosen so that $\mu\in\mathcal{G}$ and $\mu=\nu$ a.s. (see [4, Lemma 2.1]).
A sequence $\xi=\{\xi_{i}:i\in\mathbb{N}\}$ of $S$-valued random variables is exchangeable if
[…]
whenever $k_{1},\ldots,k_{m}$ are distinct elements of $\mathbb{N}$. A function $\sigma:\mathbb{N}\to\mathbb{N}$ is called a (finite) permuatation if $\sigma$ is bijective and there exists $n_{0}\in\mathbb{N}$ such that $\sigma(n)=n$ for all $n\geq n_{0}$. The sequence $\xi$ is exchangeable if and only if $\{\xi_{\sigma(i)}\}$ and $\{\xi_{i}\}$ have the same finite-dimensional distributions whenever $\sigma$ is a permutation.
By [8, Theorem 1.1], we have that $\xi$ is conditionally i.i.d. if and only if $\xi$ is exchangeable. In that case, the random measure $\mu$ in (2.1) is a.s. unique, $\sigma(\xi)$-measurable, and satisfies $\mu_{n}(B)\to\mu(B)$ a.s., for all $B\in\mathcal{S}$, where
$\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{\xi_{i}}$
are the empirical measures (see [8, Proposition 1.4]). In fact, we have that $\mu_{n}\to\mu$ a.s. in $M_{1}$ under the Prohorov metric (see [4, Proposition 2.4]). In [4, Proposition 2.5], the following conditional version of the law of large numbers is presented.
Proposition 2.1.
Let $(T,\mathcal{T})$ be a measurable space. Suppose $Y$ is a $T$-valued random variable, $\xi$ is conditionally i.i.d. given $\mathcal{G}$, and $f:T\times S\to\mathbb{R}$ is a measurable function such that $E|f(Y,\xi_{1})|<\infty$. If $Y\in\mathcal{G}$ and $Z$ is any version of $E[f(Y,\xi_{1})\mid\mathcal{G}]$ then
[…]
In particular,
[…]
2.4 Bayesian inference for row exchangeable arrays
The following four theorems come from [4]. The first is a de Finetti theorem for row exchangeability ([4, Theorem 3.2]). The final three ([4, Theorems 3.3, 3.4, 3.5]) provide needed tools for performing Bayesian inference on row exchangeable arrays.
Theorem 2.2.
Suppose $\xi$ is a row exchangeable, infinite array of $S$-valued random variables. Then there exists an a.s. unique random probability measure $\varpi$ on $M_{1}$ and an a.s. unique sequence $\mu=\{\mu_{i}:i\in\mathbb{N}\}$ of random probability measures on $S$ such that ${\mu\mid\varpi}\sim\varpi^{\infty}$ and ${\xi_{i}\mid\mu_{i}}\sim\mu_{i}^{\infty}$ for all $i\in\mathbb{N}$.
Theorem 2.3.
If $\xi$ is row exchangeable, then
[…]
for all $m,n\in\mathbb{N}$.
Theorem 2.4.
Let $\xi$ be row exchangeable. Fix $M\in\mathbb{N}$. For each $i\in\{1,\ldots,M\}$, let $N_{i},O_{i}\in\mathbb{N}$ with $N_{i}<O_{i}$, and let $A_{ij}\in\mathcal{S}$ for $i\leq M$ and $N_{i}<j\leq O_{i}$. Let $\mathcal{G}$ be a sub-$\sigma$-algebra of $\sigma(X_{1N_{1}},X_{2N_{2}},\ldots,X_{MN_{M}})$. Then
$P\left(\bigcap_{i=1}^{M}\bigcap_{j=N_{i}+1}^{O_{i}}\{\xi_{ij}\in A_{ij}\}\;\middle|\;\mathcal{G}\right)=E\left[\prod_{i=1}^{M}\prod_{j=N_{i}+1}^{O_{i}}\mu_{i}(A_{ij})\;\middle|\;\mathcal{G}\right].$
Theorem 2.5.
Let $\xi$ be row exchangeable. Fix $m,M,N\in\mathbb{N}$ with $m<M$. Then ${\bm{X}}_{mN}$ and $\{(\mu_{j},X_{jN})\}_{j=m+1}^{M}$ are conditionally independent given $\bm{\mu}_{m}$.
According to Theorem 2.4 with $N_{i}=N$, $O_{i}=N^{\prime}$, and $\mathcal{G}=\sigma({\bm{X}}_{MN})$, we can compute the posterior distribution (1.2), provided we can compute $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$. Note that
Theorem 2.5 shows that in (2.3), the first $m-1$ rows of ${\bm{X}}_{MN}$ can be omitted. Hence, the conditional distribution (1.2) is entirely determined by the conditional distributions given in (1.3).
3 Overview of the Dirichlet process
3.1 Definition
The gamma distribution on $\mathbb{R}$ is defined by
$\operatorname{Gamma}(\alpha,\lambda)(dx)\propto 1_{(0,\infty)}x^{\alpha-1}e^{\lambda x}\,dx$
for $\alpha>0$ and $\lambda>0$, and $\operatorname{Gamma}(0,\lambda)=\delta_{0}$ for $\lambda>0$. The beta distribution on $\mathbb{R}$ is defined by
$\operatorname{Beta}(\alpha,\beta)(dx)\propto 1_{(0,1)}x^{\alpha-1}(1-x)^{\beta-1}\,dx$
for $\alpha>0$ and $\beta>0$, $\operatorname{Beta}(0,\beta)=\delta_{0}$ for $\beta>0$, and $\operatorname{Beta}(\alpha,0)=\delta_{1}$ for $\alpha>0$.
Let $\alpha=(\alpha_{0},\ldots,\alpha_{d})\in[0,\infty)^{d+1}$ with $\alpha_{j}>0$ for some $j$. Let $Z_{0},\ldots,Z_{d}$ be independent with $Z_{j}\sim\operatorname{Gamma}(\alpha_{j},1)$, and define $Y_{j}=Z_{j}/(\sum_{i=1}^{k}Z_{i})$. Then the Dirichlet distribution with parameter $\alpha$ is defined to be the distribution of $(Y_{0},\ldots,Y_{d})$. We denote it by $\operatorname{Dir}(\alpha)$. It is a probability measure on the simplex, $\Delta^{d}$, defined in (1.1).
If $\alpha_{j}=0$, then $Y_{j}=0$ a.s. If $\alpha_{j}>0$ for all $j$, then $\operatorname{Dir}(\alpha)$ is the probability measure on $\Delta^{d}$ proportional to $t_{0}^{\alpha_{0}-1}\cdots t_{d}^{\alpha_{d}-1}\,dt$. Here, $dt$ denotes the surface measure on $\Delta^{d}$. For fixed $j$, we have $Y_{j}\sim\operatorname{Beta}(\alpha_{j},\sum_{i\neq j}\alpha_{i})$.
@@ 3.1 Definition
Given a nonzero, finite measure $\alpha$ on $S$, we define a Dirichlet process on $S$ with parameter $\alpha$ to be a random probability measure $\lambda$ on $S$ that satisfies
@@
whenever $\{B_{0},\ldots,B_{d}\}\subseteq\mathcal{S}$ is a partition of $S$. Let $\mathcal{D}(\alpha)$ denote the law of a Dirichlet process with parameter $\alpha$. Since a Dirichlet process is an $M_{1}$-valued random variable, it follows that $\mathcal{D}(\alpha)$ is a probability measure on $(M_{1},\mathcal{M}_{1})$. That is, $\mathcal{D}(\alpha)\in M_{1}(M_{1})$. If $B\in\mathcal{M}_{1}$, then we write $\mathcal{D}(\alpha,B)$ for $(\mathcal{D}(\alpha))(B)$.
Given $\alpha$ as above, let $\kappa=\alpha(S)>0$ and $\rho=\kappa^{-1}\alpha$, so that $\rho\in M_{1}$. We will typically write $\mathcal{D}(\alpha)=\mathcal{D}(\kappa\rho)$, and think of the law of a Dirichlet process as being determined by two parameters, a positive number $\kappa\in(0,\infty)$ and a probability measure $\rho\in M_{1}$. The measure $\rho$ is called the base measure, or base distribution, and the number $\kappa$ is called the concentration parameter. The following theorem shows that $\rho$ is the mean measure of $\lambda$ and can be found in [5, Theorem 4.3].
Theorem 3.1.
Let $\varphi:S\to\mathbb{R}$ be measurable. If $\varphi\in L^{1}(\rho)$, then
[…]
Taking $\varphi=1_{A}$ gives the special case, $E[\lambda(A)]=\rho(A)$ for all $A\in\mathcal{S}$.
Written analytically, (3.2) takes the form
$\int_{M_{1}}\int_{S}\varphi(x)\,\nu(dx)\,\mathcal{D}(\kappa\rho,d\nu)=\int_{S}\varphi(x)\,\rho(dx).$
Taking $\varphi=1_{A}$, we have
$\int_{M_{1}}\nu(A)\,\mathcal{D}(\kappa\rho,d\nu)=\rho(A),$ (3.3)
for all $A\in\mathcal{S}$.
3.2 The stick-breaking construction
The Dirichlet process can be built explicitly through the Sethuraman stick-breaking construction, which we outline below. See [14] for further details.
Let $V=\{V_{k}\}_{k=1}^{\infty}$ be i.i.d. $(0,1)$-valued random variables with $V_{k}\sim\operatorname{Beta}(1,\kappa)$. (If $\kappa=1$, then each $V_{k}$ is uniformly distributed. The larger $\kappa$ is, the smaller the $V_{k}$’s will tend to be; and conversely, the smaller $\kappa$ is, the larger the $V_{k}$’s will tend to be.) Note that by the second Borel-Cantelli lemma, $\sum_{k}V_{k}=\infty$ a.s. Choose $A\in\mathcal{F}$ such that $P(A)=1$ and $\sum_{k}V_{k}(\omega)=\infty$ for all $\omega\in A$.
We define a sequence $R=\{R_{k}\}_{k=1}^{\infty}$ of random variables as follows. On $A^{c}$, let $R_{1}=1$ and $R_{k}=0$ for $k>1$. On $A$, recursively define $R_{1}=V_{1}$ and
$R_{k+1}=\left(1-\sum_{j=1}^{k}R_{j}\right)V_{k+1}.$
The $R_{k}$’s can be thought of as the lengths of pieces of a stick of length 1. Initially, we break off a piece of length $V_{1}$. After we have broken off $n$ pieces, we take what remains and break off a proportion of it, where $V_{k+1}$ is that proportion. On $A$, the $R_{k}$’s can be expressed explicitly as
$R_{k}=V_{k}\prod_{j=1}^{k-1}(1-V_{j}).$
Note that $\sum_{k=1}^{\infty}R_{k}=1$. This is trivially the case on $A^{c}$. To see that it also holds on $A$, note that the above two displays give
$1-\sum_{j=1}^{k}R_{j}=\frac{R_{k+1}}{V_{k+1}}=\prod_{j=1}^{k}(1-V_{j}).$
Using results about infinite products, this shows that $0<\sum_{j=1}^{\infty}R_{j}\leq 1$, and $\sum_{j=1}^{\infty}R_{j}=1$ if and only if $\sum_{j=1}^{\infty}V_{j}=\infty$.
Let $\{U_{k}\}_{k=1}^{\infty}$ be i.i.d. $S$-valued random variables with $U_{k}\sim\rho$, and define
$\lambda=\sum_{k=1}^{\infty}R_{k}\delta_{U_{k}},$
so that $\lambda:\Omega\to M_{1}$ is a random probability measure that satisfies (3.1).
Note that if $\kappa$ is small, then the maximum of the $R_{k}$’s will tend to be large, meaning that $\lambda$ is likely to be concentrated on only a few points in $S$. Conversely, if $\kappa$ is large, then the $R_{k}$’s will be small, so that the mass of $\lambda$ is spread out over many points in $S$.
It can be shown, using the stick-breaking construction, that the function $(\kappa,\rho)\mapsto\mathcal{D}(\kappa\rho)$ is continuous and therefore a kernel from $(0,\infty)\times M_{1}$ to $M_{1}$. Moreover, if $\rho$ is not a point mass, then $\mathcal{D}(\kappa\rho)$ has no discrete component. That is, $\mathcal{D}(\kappa\rho,\{\nu\})=0$ for all $\nu\in M_{1}$.
3.3 Samples from a Dirichlet process
@@
$P(\eta_{i}\in A)=E[P(\eta_{i}\in A\mid\lambda)]=E[\lambda(A)]=\rho(A).$ (3.4)
Thus, $\rho$ represents our prior distribution on the individual $\eta_{i}$’s, in the case that we have not observed any of their values. The posterior distribution is given in the following theorem, which can be found in [5, Theorem 3.1].
Theorem 3.2.
With notation as above, we have
@@
for every $A\in\mathcal{S}^{n}$ and every $B\in\mathcal{M}_{1}$.
@@
Let $\lambda$ and $\eta$ be as in Theorem 3.2. Then
@@
[…]
@@
[…]
3.4 Mixtures and noisy observations
[…]
Let $\alpha$ be a nonzero, finite random measure on $S$. That is, $\alpha$ is a random measure on $S$ such that $\alpha(S)\in(0,\infty)$ a.s. Let $\lambda$ be a random probability measure on $S$ such that $\lambda\mid\alpha\sim\mathcal{D}(\alpha)$. (This notation is meaningful, since $\alpha\mapsto\mathcal{D}(\alpha)$ is a probability kernel.) In this case we call $\lambda$ a mixture of Dirichlet processes on $S$ with mixing distribution $\mathcal{L}(\alpha)$. We also let $\kappa=\alpha(S)$ and $\rho=\alpha/\alpha(S)$, so that $\kappa$ and $\rho$ are random variables taking values in $(0,\infty)$ and $M_{1}$, respectively.
Let $\lambda$ be a mixture of Dirichlet processes as above. A sequence of samples from $\lambda$ is a sequence $\eta=\{\eta_{i}\}_{i=1}^{\infty}$ that satisfies ${\eta\mid\lambda,\alpha}\sim\lambda^{\infty}$. The following proposition generalizes (3.5) to mixtures.
Proposition 3.4.
With notation as above, we have
$\mathcal{L}(\lambda\mid\bm{\eta}_{n},\alpha)=\mathcal{D}\bigg{(}\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}}\bigg{)},$

### v2

@@ 1.3.3 The infinite relational model
    the entries of the array $\{\xi_{ij}\}$ are conditionally i.i.d. given $\mu$,
whereas a separately exchangeable array does not. Because of (iii), our process satisfies $\mathcal{L}({\bm{X}}_{mn}\mid\mu)=\prod_{i=1}^{m}\mu_{i}^{n}$, a fact that we will use again in Section 4.2. In particular, we have $\mathcal{L}(\xi_{11},\xi_{21}\mid\mu_{1},\mu_{2})=\mu_{1}\times\mu_{2}$, and this is true regardless of whether $\mu_{1}=\mu_{2}$ or not.
Now, suppose we modify our process so that whenever $\mu_{i}=\mu_{i^{\prime}}$, the rows $\xi_{i}$ and $\xi_{i^{\prime}}$ are identical. That is, $\mu_{i}=\mu_{i^{\prime}}$ implies $\xi_{ij}=\xi_{i^{\prime}j}$ for all $j$. In this case, the array $\xi$ no longer satisfies (iii) and is no longer row exchangeable. It is, however, separately exchangeable. In fact, the array $\xi$, modified in this way, would be an instance of an IRM called a simple IRM. A general IRM is obtained from a simple IRM through a process called randomization.
We can apply randomization to any random array, so let use drop any specific assumptions for now and just let $\xi$ be an arbitrary random array of elements of $S$. To obtain a randomization of $\xi$, let $(T,\mathcal{T})$ be a measurable space and let $Q$ be a probability kernel from $S$ to $T$. Define $\Xi=\{\Xi_{ij}\}$ so that $\{\Xi_{ij}\}$ is conditionally i.i.d. given $\xi$ and $\Xi_{ij}\mid\xi\sim Q(\xi_{ij})$. Then the array $\Xi$ is called a $Q$-randomization of $\xi$. If we apply $Q$-randomization to the simple IRM in the previous paragraph, then we obtain a general IRM.
The IRM is a model that is used when the columns of $\xi$ correspond to fundamentally distinct objects that remain fixed from row to row. For instance, suppose the agents are movie critics, and their $j$th action is to review the $j$th movie on some fixed list of movies that is shared by all critics. Suppose we knew the exact reviewing tendencies of the first two critics. That is, we know $\mu_{1}$ and $\mu_{2}$. Even then, if we observe $\xi_{11}$, the first critic’s review of the first movie, then we would learn something about the first movie. This could potentially affect our probabilities for $\xi_{21}$, the second critic’s review of the first movie. In other words, $\xi_{11}$ and $\xi_{21}$ would not be conditionally independent given $\mu_{1}$ and $\mu_{2}$. This is distinctly different from the motivating example in [17], where the agents are different medical centers, and the actions are the individual patient outcomes produced by those centers. In the movie critic scenario, an NDP is not an appropriate model, whereas an IRM would be reasonable. In the language of the machine learning literature, the rows, which represents the agents, exhibit clustering in both scenarios. But only in the movie critic scenario do the columns exhibit clustering.
1.4 Outline of paper
In Section 2, we give some necessary background information and establish the notational conventions that we will use throughout the paper. In Section 3, we give a precise formulation of the method of sequential imputation as it appears in [10]. This formulation is given in Theorem 3.4. Liu’s error in [11] is to apply this result to the simple NDP with $S=\{0,1\}$ despite the fact that the hypotheses do not hold. To correct this error, we give a new proof under weaker hypotheses, and present this in Theorem 3.6.
In Section 4, we apply Theorem 3.6 to the NDP with a general state space $S$. Our main results are Theorem 4.1 and Corollary 4.2. Finally, in Section 5, we present several hypothetical examples to illustrate the use of our main results. See https://github.com/jason-swanson/ndp for the code used to generate the simulations in Section 5.
2 Notation and background
2.1 General notation
Throughout the paper, we fix a complete and separable metric space $S$. When needed, we let $\mathcal{S}$ denote its Borel $\sigma$-algebra. As noted in the introduction, we write $M_{1}=M_{1}(S)$ for the set of Borel probability measures on $S$, and equip $M_{1}$ with the Prohorov metric so that $M_{1}$ is itself a complete and separable metric space.
Let $(T,\mathcal{T})$ be a measurable space. Recall that a probability kernel from $T$ to $S$ is a measurable function $\mu:T\to M_{1}(S)$. If $\mu$ is any function from $T$ to $M_{1}(S)$, measurable or not, we write $\mu(t,B)$ for $(\mu(t))(B)$. Such a function is a kernel if and only if $\mu(\cdot,B)$ is measurable for each Borel set $B$. Note that a random probability measure on $S$ is a probability kernel from $\Omega$ to $S$. Also, if $\mu$ is a probability kernel from $T$ to $S$ and $Y$ is a $T$-valued random variables, then $\mu(Y)$ is a random probability measure on $S$.
@@ 2.1 General notation
Now let $S^{\prime}$ be another complete and separable metric space. Let $\gamma$ be a probability kernel from $T$ to $S$ and $\gamma^{\prime}$ a probability kernel from $T\times S$ to $S^{\prime}$. (We allow the possibility that $T$ is a singleton, in which case $\gamma$ is a probability measure on $S$ and $\gamma^{\prime}$ is a probability kernel from $S$ to $S^{\prime}$.) We write $\gamma\gamma^{\prime}$ to denote the probability kernel from $T$ to $S\times S^{\prime}$ characterized by
$(\gamma\gamma^{\prime})(y,A\times A^{\prime})=\int_{A}\gamma^{\prime}(y,z,A^{\prime})\,\gamma(y,dz).$
In particular, this means
@@
We write $\mathcal{L}(X)$ and $\mathcal{L}(X\mid Y)$ for the distribution of $X$ and the regular conditional distribution of $X$ given $Y$, respectively. We use semicolons to indicate evaluation, so that $\mathcal{L}(X;B)=(\mathcal{L}(X))(B)$ and $\mathcal{L}(X\mid Y;B)=P(X\in B\mid Y)$. We also adopt the usual notation, $X\sim\mu$ and $X\mid Y\sim\mu$, to mean $\mathcal{L}(X)=\mu$ and $\mathcal{L}(X\mid Y)=\mu(Y)$, respectively. In the case $\mathcal{L}(X\mid Y)=\mu(Y)$, the probability kernel $\mu$ is only determined $\mu(Y)$-a.e. Nonetheless, if a particular $\mu$ has been fixed, we use the notation $\mathcal{L}(X\mid Y=y)$ to denote the probability measure $\mu(y,\cdot)$.
2.2 Dirichlet processes
@@
Given a nonzero, finite measure $\alpha$ on $S$, a Dirichlet process on $S$ with parameter $\alpha$ is a random probability measure $\lambda$ on $S$ that satisfies
@@
whenever $\{B_{0},\ldots,B_{d}\}\subseteq\mathcal{S}$ is a partition of $S$. The right-hand side of (2.1) is the Dirichlet distribution on the simplex $\Delta^{d}$. We write $\mathcal{D}(\alpha)$ to denote the law of a Dirichlet process with parameter $\alpha$. Since a Dirichlet process is an $M_{1}$-valued random variable, it follows that $\mathcal{D}(\alpha)$ is a Borel probability measure on $M_{1}$. That is, $\mathcal{D}(\alpha)\in M_{1}(M_{1})$. Given a Borel set $B\subseteq M_{1}$, we write $\mathcal{D}(\alpha,B)$ for $(\mathcal{D}(\alpha))(B)$.
With $\alpha$ as above, let $\kappa=\alpha(S)>0$ and $\rho=\kappa^{-1}\alpha$, so that $\rho\in M_{1}$. We typically write $\mathcal{D}(\alpha)=\mathcal{D}(\kappa\rho)$, and think of the law of a Dirichlet process as being determined by two parameters, a positive number $\kappa\in(0,\infty)$ and a probability measure $\rho\in M_{1}$. We call the measure $\rho$ the base measure, or base distribution, and the number $\kappa$ the concentration parameter. If $\varphi\in L^{1}(\rho)$, then
$E\int\varphi\,d\lambda=\int\varphi\,d\rho.$
Taking $\varphi=1_{A}$ gives the special case, $E[\lambda(A)]=\rho(A)$. This and other basic properties of the Dirichlet process can be found in [6].
@@
$P(\eta_{i}\in A)=E[P(\eta_{i}\in A\mid\lambda)]=E[\lambda(A)]=\rho(A).$
Thus, $\rho$ represents our prior distribution on the individual $\eta_{i}$, in the case that we have not observed any of their values. As shown in [6, Theorem 3.1], the posterior distribution is given by
@@
for every $A\in\mathcal{S}^{n}$ and every Borel set $B\subseteq M_{1}$.
@@
We first note that
@@
$\int_{S^{n+1}}f\,d\rho_{n+1}=\frac{\kappa}{\kappa+n}E\bigg[\int_{S}{f(\bm{\eta}_{n},x_{n+1})}\,\rho(dx_{n+1})\bigg]+\frac{1}{\kappa+n}\sum_{i=1}^{n}E[f(\bm{\eta}_{n},\eta_{i})].$
@@
$\int_{S^{n+1}}f\,d\rho_{n+1}=\frac{\kappa}{\kappa+n}\,{\int_{S^{n}}\int_{S}f({\bm{x}}_{n+1})\,\rho(dx_{n+1})\,\rho_{n}(d{\bm{x}}_{n})}\\ +\frac{1}{\kappa+n}\,{\sum_{i=1}^{n}\int_{S^{n}}f({\bm{x}}_{n},x_{i})\,\rho_{n}(d{\bm{x}}_{n})},$ (2.5)
for all $n\in\mathbb{N}$.
2.3 Mixtures of Dirichlet processes
Now let $\alpha$ be a random measure on $S$ such that $\alpha(S)\in(0,\infty)$ a.s. Let $\lambda$ be a random probability measure on $S$ such that $\lambda\mid\alpha\sim\mathcal{D}(\alpha)$. In this case, $\lambda$ is called a mixture of Dirichlet processes on $S$ with mixing distribution $\mathcal{L}(\alpha)$. We also let $\kappa=\alpha(S)$ and $\rho=\alpha/\alpha(S)$, so that $\kappa$ and $\rho$ are random variables taking values in $(0,\infty)$ and $M_{1}$, respectively. Some fundamental properties of these mixtures are given in [1].
A sequence of samples from $\lambda$ is a sequence $\eta=\{\eta_{i}\}_{i=1}^{\infty}$ that satisfies ${\eta\mid\lambda,\alpha}\sim\lambda^{\infty}$. In this case, (2.2) generalizes to
$\mathcal{L}(\lambda\mid\bm{\eta}_{n},\alpha)=\mathcal{D}\bigg(\alpha+\sum_{i=1}^{n}\delta_{\eta_{i}}\bigg),$ (2.6)